### v2

# Parametric Channel Estimation and Design for Active-RIS-Assisted Communications

Abstract
Reconfigurable Intelligent Surface (RIS) technology has emerged as a key enabler for future wireless communications. However, its potential is constrained by the difficulty of acquiring accurate user-to-RIS channel state information (CSI), due to the cascaded channel structure and the high pilot overhead of non-parametric methods. Unlike a passive RIS, where the reflected signal suffers from multiplicative path loss, an active RIS amplifies the signal, improving its practicality in real deployments. In this letter, we propose a parametric channel estimation method tailored for active RISs. The proposed approach integrates an active RIS model with an adaptive Maximum Likelihood Estimator (MLE) to recover the main channel parameters using a minimal number of pilots. To further enhance performance, an adaptive active RIS configuration strategy is employed, which refines the beam direction based on an initial user location estimate. Moreover, an orthogonal angle-pair codebook is used instead of the conventional Discrete Fourier Transform (DFT) codebook, significantly reducing the codebook size and ensuring reliable operation for both far-field and near-field users. Extensive simulations demonstrate that the proposed method achieves near-optimal performance with very few pilots compared to non-parametric approaches. Its performance is also benchmarked against that of a traditional passive RIS under the same total power budget to ensure fairness. Results show that active RIS yields higher spectral efficiency (SE) by eliminating the multiplicative fading inherent in passive RISs and allocating more resources to data transmission.
Index Terms:
Active Reconfigurable Intelligent Surface, Channel Estimation, MLE.
I Introduction
Next-generation (6G) wireless networks require mechanisms that reconfigure the propagation environment to improve end-to-end performance. Reconfigurable intelligent surfaces (RISs) address this need by enhancing wireless channels via programmable metasurfaces [15]. A passive RIS uses many low-cost scatterers to reflect incident signals with adjustable phase shifts, shaping the channel without extra transmit power [6]. However, passive RIS suffers from a double-fading effect [9], motivating architectures that can compensate for the severe cascaded path loss. To overcome this limitation, active RIS (ARIS) architectures integrate active components that amplify rather than merely reflect the impinging wave.
Even with these advantages, reliably acquiring channel state information (CSI) in RIS-aided systems remains challenging. Conventional schemes require the Base Station/User to transmit pilots while the RIS cycles through many configurations, so the pilot typically scales with the number of elements, which becomes prohibitive for large surfaces and under mobility [13]. Mobility-aware designs based on statistical temporal correlation can track small-scale fading in localized regions, but they assume stationary statistics and still require pilot lengths on the order of the RIS size, making them unsuitable for non-stationary motion or large coverage areas [14]. Subarray-based reduced subspace methods can shrink the effective training dimension, yet either sacrifice beamforming gain or still incur substantial overhead for physically large RISs [4]. Parametric approaches exploit the low-rank of mmWave channels to reduce cascaded-channel estimation into angle and gain estimation problems using matrix factorization, compressive sensing, or atomic norm minimization but are mainly developed for line-of-sight (LoS) dominant far-field scenarios [8]. Recent near-field works for XL-MIMO and RIS for THz systems introduce distance aware parametric models and polar-domain sparsity. However, they are still in an early stage and often treat only the RIS path [3].
Building on these observations, Haghshenas et al. proposed in [7] a parametric MLE framework that directly targets most of these issues for passive RISs. Their method jointly estimates the direct BS–user channel and the LoS user–RIS channel, recovering angular parameters in the far field and, in the near field, the user–RIS distance, so that both near and far-field regimes are handled within a single parametric model. To curb pilot overhead, they combine this MLE with an adaptive RIS configuration strategy, and a wide-beam initialization further ensures useful observations even at low SNR and under mobility. The framework is, however, developed under the standard ideal passive RIS assumptions: each RIS element is modeled as a lossless phase shifter with perfect control, and the signal model includes only the usual receiver thermal noise, which is appropriate for a purely passive surface. Even with such improved channel acquisition, a passive RIS link remains fundamentally limited by the multiplicative fading effect. To overcome this limitation, one needs active surfaces that compensate for it by amplifying the reflected signal.
Very recently, researchers have shifted their focus to active RIS-aided CSI. Authors in [11] applied MLE to an active RIS-aided LoS channel for user localization, showing that parametric modeling can achieve high accuracy with active surfaces. The authors in [2] analyzed hardware impairment effects by deriving the Cramér–Rao bound for active RIS systems and proposing a joint least-squares channel estimator. In both these cases, the primary focus was the mean-squared error (MSE) rather than pilot overhead. In a complementary direction, Lee et al. [10] studied a semi-passive RIS with a few active sensing elements and reconstructed the full RIS–user channel via a correlation-aware linear combination of partial CSI, so that the training overhead no longer scales with the total number of RIS elements. However, neither does it exploit the per-element amplification nor the gain and phase calibration constraints.
Motivated by these gaps, we propose a parametric channel estimation methodology tailored to active RIS architectures. Specifically, we formulate a maximum likelihood estimator that yields closed-form estimates for all relevant channel parameters under both near and far-field conditions. The estimator jointly recovers the full complex gain (amplitude and phase) of the channel and explicitly obtains the user’s geometric parameters: angles of arrival and departure for far-field users and angles plus distance for near-field users within a unified framework. We further integrate an adaptive training and beamforming strategy that iteratively refines the RIS configuration based on initial estimates, concentrating pilot power in the most informative directions. This active RIS design aligns both amplification and phase shifts with the parametric estimation objective, thereby improving accuracy without increasing pilot length. To the best of our knowledge, this is the first comprehensive parametric channel estimator developed specifically for active RIS hardware, accounting for amplified noise and operating seamlessly across near and far-field regimes.
II System and Channel Model
We consider an uplink system comprising a single-antenna base station (BS) and a single-antenna user equipment (UE) assisted by an active reconfigurable intelligent surface (ARIS). No direct BS-to-UE link is assumed. The active RIS comprises $N$ elements with independently tunable amplitude and phase. We model the BS-to-ARIS channel, $\bm{h}\in\mathbb{C}^{N}$, as a deterministic LoS link. Given that the BS and active RIS are fixed in location, this channel is considered quasi-stationary and non-fading over the signaling interval. The ARIS-to-UE channel is $\bm{g}\triangleq[g_{1},\ldots,g_{N}]^{\mathrm{T}}\in\mathbb{C}^{N}$, which is time varying due to user mobility. To maintain a high signal-to-noise ratio (SNR), the BS periodically estimates the channel and updates the active RIS coefficients (phase shifts and gains) accordingly.
Let the UE transmit the symbol $x\sim\mathcal{CN}(0,P_{d})$. The received signal at the BS is:
$y\;=\;\big(\bm{\Phi}^{\mathrm{T}}\mathbf{D}_{h}\bm{g}\big)x\;+\;\bm{\Phi}^{\mathrm{T}}\mathbf{D}_{h}\bm{v}\;+\;w,$ (1)
where $\bm{\Phi}\triangleq[\Phi_{1},\ldots,\Phi_{N}]^{\mathrm{T}}$ , it collects the complex active RIS coefficients. Here $\mathbf{D}_{h}=\operatorname{diag}(h_{1},\ldots,h_{N})$, and $\Phi_{n}=p_{n}e^{\mathrm{j}\theta_{n}}$ is the active RIS configuration, where $\theta_{n}$ is the phase shift and $p_{n}$ is the amplification factor of the $n$th element, respectively.
Two noise components are present: the active RIS amplification noise, $\bm{v}\sim\mathcal{CN}(\mathbf{0}_{N},\,\sigma_{v}^{2}\mathbf{I}_{N})$, and the receiver noise, $w\sim\mathcal{CN}(0,\,\sigma^{2})$.
The spectral efficiency (SE) can be expressed as:
$R=\log_{2}\!\left(1+\frac{P_{d}\left|\bm{\Phi}^{\mathrm{T}}\mathbf{D}_{h}\bm{g}\right|^{2}}{\sigma^{2}+\left\|\bm{\Phi}^{\mathrm{T}}\mathbf{D}_{h}\right\|_{2}^{2}\sigma_{v}^{2}}\right).$ (2)
The LoS channel vector from UE-to-ARIS is modeled as:
$\bm{g}\;=\;\sqrt{\beta}\,e^{\mathrm{j}\omega}\,\bm{a}(\psi),$ (3)
where $\beta\geq 0$ and $\omega\in[0,2\pi)$ denote, respectively, the channel gain and the phase referenced to the first active RIS element, and $\bm{{a}(\psi)}$ is the array-response vector parameterized by $\bm{\psi}$. This model applies to both near and far-field propagation: in the far field, $\bm{\psi}=(\varphi,\phi)$ comprises the azimuth $\varphi$ and elevation $\phi$; in the near field, $\bm{\psi}=(\varphi,\phi,r)$ additionally includes the distance $r$ from the UE to the first active RIS element (Reference element).
For the near field, the array response vector can be written as:
$\mathbf{a}(\varphi,\phi,r)=\begin{bmatrix}1,\,e^{-\mathrm{j}\frac{2\pi}{\lambda}s_{2}},\,e^{-\mathrm{j}\frac{2\pi}{\lambda}s_{3}},\,\ldots,\,e^{-\mathrm{j}\frac{2\pi}{\lambda}s_{N}}\end{bmatrix}^{\mathrm{T}},$ (4)
where $s_{n}\triangleq r_{n}-r_{1}$ denotes the path-length difference between the $n$th element and the reference element, with $r_{n}$ being the distance from the UE to the $n$th active RIS element.
In the far field, the response reduces to the conventional planar-wave form [1]
$a(\varphi,\phi)=\bigg[$ $1,\ldots,e^{-j\frac{2\pi}{\lambda}\left[i(n)\Delta_{\mathrm{H}}\sin(\varphi)\cos(\phi)+j(n)\Delta_{\mathrm{V}}\sin(\phi)\right]},$ (5)
$\ldots,e^{-j\frac{2\pi}{\lambda}\left[(N_{\mathrm{H}}-1)\Delta_{\mathrm{H}}\sin(\varphi)\cos(\phi)+(N_{\mathrm{V}}-1)\Delta_{\mathrm{V}}\sin(\phi)\right]}\bigg]^{\mathrm{T}},$
where $i(n)$ and $j(n)$ are the horizontal and vertical coordinates of the $n$th active RIS element, $\Delta_{H}$ and $\Delta_{V}$ are the inter-element spacings along the horizontal and vertical axes, and $N_{H}$ and $N_{V}$ denote the numbers of elements per row and column, respectively. We can express $i(n)$ and $j(n)$ in terms of $\Delta_{H}$ and $\Delta_{V}$ [1], $i(n)=\operatorname{mod}(n-1,N_{\mathrm{H}})\quad\text{and}\quad j(n)=\left\lfloor\frac{n-1}{N_{\mathrm{H}}}\right\rfloor$.
III MLE-based Estimation for active RIS
We consider the received signal model when the UE transmits a pilot symbol as:
$\bm{y}\;=\;\big(\bm{B}\,\mathbf{D}_{h}\bm{g}\big)\,\sqrt{P_{p}}\;+\;\bm{B}\,\mathbf{D}_{h}\bm{v}\;+\;\bm{w},$ (6)
where $\bm{y}\in\mathbb{C}^{L}$ is the received vector across $L$ instances of the pilot, $\bm{B}\in\mathbb{C}^{L\times N}$ is the active RIS configuration matrix over these instances, $P_{p}$ is the pilot power. We are sending the same pilot but for $L$ slots.
The deterministic mean is $\bm{\mu}(\bm{g})\;=\;\big(\bm{B}\,\mathbf{D}_{h}\bm{g}\big)\,\sqrt{P_{p}},$ and the effective noise is $\tilde{\bm{w}}\;=\;\bm{B}\,\mathbf{D}_{h}\bm{v}\;+\;\bm{w};\qquad\tilde{\bm{w}}\sim\mathcal{CN}\!\left(\bm{0},\,\bm{F}\right),$ with total noise covariance,
$\bm{F}\;=\;\sigma^{2}\mathbf{I}_{L}\;+\;\sigma_{v}^{2}\,\bm{B}\,\mathbf{D}_{h}\mathbf{D}_{h}^{\mathrm{H}}\bm{B}^{\mathrm{H}}.$ (7)
The Maximum Likelihood (ML) estimate of the channel $\bm{g}$, denoted by $\hat{\bm{g}}$, is the argument that maximizes the likelihood function of the received signal vector $\bm{y}$. The likelihood function $f_{\bm{Y}}(\bm{y};\bm{g})$ is given by
$f_{\bm{Y}}(\bm{y};\bm{g})=\frac{1}{\pi^{L}\det(\bm{F})}\,\exp\left[-\big(\bm{y}-\bm{\mu}(\bm{g})\big)^{\mathrm{H}}\bm{F}^{-1}\big(\bm{y}-\bm{\mu}(\bm{g})\big)\right].$ (8)
Maximizing the likelihood $f_{\bm{Y}}(\bm{y};\bm{g})$ is equivalent to minimizing the quadratic form in the exponent. Since $\bm{F}$ is Hermitian positive definite, $\bm{F}^{-1}$ is Hermitian. Therefore,
$\big(\bm{y}-\bm{\mu}(\bm{g})\big)^{\mathrm{H}}\bm{F}^{-1}\big(\bm{y}-\bm{\mu}(\bm{g})\big)$ (9)
$=\bm{y}^{\mathrm{H}}\bm{F}^{-1}\bm{y}-2\sqrt{P_{p}}\,\Re\!\left\{\bm{y}^{\mathrm{H}}\bm{F}^{-1}\big(\bm{B}\,\mathbf{D}_{h}\bm{g}\big)\right\}$
$\quad+P_{p}\big(\bm{B}\,\mathbf{D}_{h}\bm{g}\big)^{\mathrm{H}}\bm{F}^{-1}\big(\bm{B}\,\mathbf{D}_{h}\bm{g}\big).$
The constant term $\bm{y}^{\mathrm{H}}\bm{F}^{-1}\bm{y}$ can be discarded as it does not depend on $\bm{g}$ and thus does not affect its estimation. Hence,
$\hat{\bm{g}}=\operatorname*{arg\,min}_{\bm{g}}\big[$ $-2\sqrt{P_{p}}\,\Re\!\left\{\bm{y}^{\mathrm{H}}\bm{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{g}\right\}$
$+P_{p}\,(\mathbf{B}\mathbf{D}_{h}\bm{g})^{\mathrm{H}}\bm{F}^{-1}(\mathbf{B}\mathbf{D}_{h}\bm{g})\big].$ (10)
Now substitute $\bm{g}=\sqrt{\beta}\,e^{\mathrm{j}\omega}\,\bm{a(\psi)}$:
$\{\hat{\beta},\,\hat{\omega},\,\hat{\bm{\psi}}\}$ $=\operatorname*{arg\,min}_{\begin{subarray}{c}\beta\geq 0,\,\omega\in[0,2\pi)\\ \bm{\psi}\in\Psi\end{subarray}}\Big[P_{p}\beta\,\bm{a(\psi)}^{\mathrm{H}}\mathbf{D}_{h}^{\mathrm{H}}\mathbf{B}^{\mathrm{H}}\bm{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a(\psi)}$ (11)
$\quad-2\sqrt{P_{p}\beta}\,\Re\!\left\{e^{\mathrm{j}\omega}\,\bm{y}^{\mathrm{H}}\bm{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a(\psi)}\right\}\Big].$
III-A ARIS Reflected Channel Phase $(\omega)$
The expression in (11) is minimized when the real part inside the objective is maximized. Therefore, the optimal phase $\hat{\omega}$ aligns the exponential term with the phase of the complex inner product as:
$\hat{\omega}\,=\,-\,\arg\!\left\{\mathbf{y}^{\mathrm{H}}\mathbf{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a}(\psi)\right\}.$ (12)
Now, by substituting the value of $\hat{\omega}$ into the equation (11) , we get:
$\{\hat{\beta},\,\,\hat{\bm{\psi}}\}$ $=\mathop{\arg\min}\limits_{\begin{subarray}{c}\beta\geq 0\\ \bm{\psi}\in\Psi\end{subarray}}\Bigg[P_{p}\beta\left(\bm{a(\psi)}^{H}\mathbf{D}_{h}^{H}\mathbf{B}^{H}\mathbf{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a(\psi)}\right)$ (13)
$\quad-2\sqrt{P_{p}}\sqrt{\beta}\,\left|\mathbf{y}^{H}\mathbf{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a(\psi)}\right|\Bigg].$
III-B ARIS Reflected Channel Amplitude $(\beta)$
This is a quadratic function in terms of $\sqrt{\beta}$, for which the minimizer can be determined in closed form as:
$\hat{\beta}\;=\;\frac{\left|\mathbf{y}^{\mathrm{H}}\mathbf{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a}(\psi)\right|^{2}}{P_{p}\left(\bm{a}(\psi)^{\mathrm{H}}\mathbf{D}_{h}^{\mathrm{H}}\mathbf{B}^{\mathrm{H}}\mathbf{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a}(\psi)\right)^{2}}.$ (14)
III-C ARIS Reflected Channel Array Response Parameter ($\bm{\psi}$)
By replacing this value $\hat{\beta}$ in equation (13), we get
$\hat{\bm{\psi}}=\mathop{\arg\max}\limits_{\bm{\psi}\in\Psi}\frac{\left|\mathbf{y}^{H}\mathbf{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a(\psi)}\right|^{2}}{\bm{a(\psi)}^{H}\mathbf{D}_{h}^{H}\mathbf{B}^{H}\mathbf{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a(\psi)}}.$ (15)
IV Adaptive Beam and Amplification Factor
Codebook & Search Space: Let $\Theta=\{\bm{\theta}_{k}\in\mathbb{C}^{N}:\ \|\bm{\theta}_{k}\|_{\infty}=1,\ k=1,\ldots,K\}$ collect phase-only beams generated from orthogonal angle pairs (Same codebook can be used for near field too, according to Weyl identity [12]). By construction, $\bm{\theta}_{i}^{H}\bm{\theta}_{j}\approx 0$ for $i\neq j$, so the search contracts from an $N\times N$ DFT matrix to $R\ll N^{2}$ nearly orthogonal beams, each column representing a single RIS configuration direction.
Initialization via Wide-Beam Probing: Following [7, Sec. VI], two complementary wide beams with unit-modulus phase vectors $\bm{\theta}_{w,1}$ and $\bm{\theta}_{w,2}$ is designed. The active RIS is configured with equal per-element gain and the wide-beam phases as:
$\bm{\Phi}_{1}=\sqrt{\frac{P_{\mathrm{RIS}}}{N}}\ \bm{\theta}_{w,1},\qquad\bm{\Phi}_{2}=\sqrt{\frac{P_{\mathrm{RIS}}}{N}}\ \bm{\theta}_{w,2}.$ (16)
The configuration matrix B is initialized with these two rows; each new pilot appends its active RIS configuration as an additional row. These pilots illuminate the UE in both near and far-field regimes and reduce the initial pilot count. $P_{\mathrm{RIS}}$ represents the total allowable active RIS power.
Active-RIS Amplitude Shaping & Phase Alignment With $\bm{\hat{g}}$, per-element phases for coherent combining across the cascaded links is taken by
$),\bar{\bm{\theta}}^{\mathrm{opt}}\;=\;\exp\!\big(-\,\mathrm{j}\,\angle(\mathbf{D}_{h}\widehat{\bm{g}}$$\big)$, (16)
which cancels the aggregate phase $h_{n}\widehat{g}_{n}$ at each element. Under the active-RIS power constraint using [5, Sec. V, Cor. 3], the per-element amplification can be computed as:
$p_{n}^{(\mathrm{est})}=C\,\frac{\alpha_{n}}{\beta_{n}+\gamma_{n}},$ (18)
where, $\alpha_{n}=|\widehat{g}_{n}|\,|h_{n}|,$ $\beta_{n}=|h_{n}|^{2},$ $\gamma_{n}=\tfrac{|\widehat{g}_{n}|^{2}\,\tfrac{p_{d}}{\sigma_{v}^{2}}+1}{\tfrac{P_{\mathrm{ris}}}{\sigma^{2}}},$ $C=\Big(\sum_{n=1}^{N}\frac{|\alpha_{n}|^{2}\,\gamma_{n}}{(\beta_{n}+\gamma_{n})^{2}}\Big)^{-1/2}.$
For the initial pilots, the channel is unknown, so this amplification is not applied. After estimating the channel from the initial pilots, the amplification factors are updated and transmitted through the next pilot using
$\bm{\Phi}^{\star}=\bm{p}^{(\mathrm{est})}\odot\bar{\bm{\theta}}_{\mathrm{opt}},$ (20)
where $\odot$ denotes the Hadamard product.
Adaptive Pilot Selection The estimate can be refined by choosing the next phase-only codeword that best aligns with the current active vector
$\bm{\theta}^{\text{next}}\,=\,\arg\max_{\bm{\theta}\in\bm{\Theta}}\bigl|\,(\bm{\Phi}^{\star})^{\mathrm{H}}\bm{\theta}\,\bigr|.$ (17)
and set
$\bm{\Phi}_{\mathrm{new}}=\bm{p}^{(\mathrm{est})}\odot\bm{\theta}_{\mathrm{next}}.$ (22)
We append $\bm{\Phi}_{\text{new}}$ as a new row to the configuration matrix $\mathbf{B}$, collect the corresponding measurement, and update $\widehat{\bm{g}}$ via the MLE step. This procedure is repeated greedily until the pilot budget is exhausted. Each selected codeword is removed from $\Theta$.
Complexity & Applicability Each refinement evaluates the remaining phase-only codewords $\Theta_{\mathrm{rem}}$ using $|\Theta_{\mathrm{rem}}|$ length-NN inner products, i.e., $\mathcal{O}\!\left(N|\Theta_{\mathrm{rem}}|\right)$ per iteration, for a total training complexity $\mathcal{O}(LNR)$ with $L$ pilots and codebook size $R$ ($R\ll N^{2}$), greatly reducing cost versus DFT exhaustive search.
Algorithm 1 Parametric MLE for Near/Far Field with Active RIS and Adaptive Pilots
1:codebook $\Theta$ (phase-only), initial wide beams $\{\bm{\theta}_{w,1},\bm{\theta}_{w,2}\}$, slots $L$, powers $(P_{p},P_{d},P_{\rm RIS})$, $\mathbf{D}_{h}=\mathrm{diag}(\mathbf{h})$
[…]
3:for $\ell=2$ to $L$ do
4:  Measurement: $\mathbf{y}_{\ell}\leftarrow\sqrt{P_{p}}\,\mathbf{B}_{\ell}\mathbf{D}_{h}\mathbf{g}+\mathbf{v}_{\ell}+\mathbf{w}_{\ell}$
5:  where $\mathbf{v}_{\ell}(i)=\mathbf{B}_{\ell}(i,:)\mathbf{D}_{h}\mathbf{v}_{i}$
6:  Parametric MLE:
7:  $\hat{\bm{\psi}}_{\ell}=\arg\max_{\bm{\psi}\in\Psi}\frac{\big|\mathbf{y}_{\ell}^{H}\mathbf{F}^{-1}\mathbf{B}_{\ell}\mathbf{D}_{h}\bm{a}(\bm{\psi})\big|^{2}}{\bm{a}(\bm{\psi})^{H}\mathbf{D}_{h}^{H}\mathbf{B}_{\ell}^{H}\mathbf{F}^{-1}\mathbf{B}_{\ell}\mathbf{D}_{h}\bm{a}(\bm{\psi})}$
8:  Closed-form updates: compute $\hat{\beta}_{\ell}$, $\hat{\omega}_{\ell}$ and set $\hat{\mathbf{g}}_{\ell}$
9:  Active amplitude shaping:
10:  for $k=1$ to NN do
11:   $p_{k}^{(\mathrm{est})}\leftarrow C\,\alpha_{k}/(\beta_{k}+\gamma_{k})$
12:  end for
13:  if $\ell<L$ then
14:   Phase steering:
15:   $)\displaystyle\bar{\bm{\theta}}_{\ell}^{\rm opt}\leftarrow\exp\!\big(-j\,\angle(\mathbf{D}_{h}\hat{\mathbf{g}}_{\ell}$$\big)$,
16:   $\bm{\Phi}_{\ell}^{\star}\leftarrow\mathbf{p}^{(\mathrm{est})}\odot\bar{\bm{\theta}}_{\ell}^{\rm opt}$.
17:   Pilot selection: $\bm{\theta}_{\ell+1}\leftarrow\arg\max_{\bm{\theta}\in\bm{\Theta}}\big|(\bm{\Phi}_{\ell}^{\star})^{H}\bm{\theta}\big|$
18:   Append & update:
19:   $\bm{\Phi}_{\rm new}\leftarrow\mathbf{p}^{(\mathrm{est})}\odot\bm{\theta}_{\ell+1}$,
20:   $\mathbf{B}_{\ell+1}\leftarrow\begin{bmatrix}\mathbf{B}_{\ell}\\ \bm{\Phi}_{\rm new}^{\top}\end{bmatrix}$,
21:   $\bm{\Theta}\leftarrow\bm{\Theta}\setminus\{\bm{\theta}_{\ell+1}\}$.
22:   Transmit pilot $\ell{+}1$; acquire $\mathbf{y}_{\ell+1}$
23:  end if
24:end for
25:Output: $\hat{\mathbf{g}}\leftarrow\hat{\mathbf{g}}_{L}$
Figure 1: NMSE versus the pilot length when the user is at a random location in the near-field of the active RIS.
Figure 2: The average Achievable Rate versus the pilot length when the user is at a random location in the near-field of the active RIS.
Figure 3: NMSE versus the pilot length when the user is at a random location in the far-field of the active RIS.
Figure 4: The average Achievable Rate versus the pilot length when the user is at a random location in the far-field of the active RIS.
Figure 5: Achievable Rate versus the percentage of total power given to active RIS for three different total powers.
V Simulation Results Analysis
We evaluate the proposed estimator via Monte Carlo simulations at a carrier of $28$ GHz with $1$ MHz bandwidth. The system has a single-antenna BS & UE assisted by an active RIS (ARIS): a $32\times 32$ UPA with half-wavelength spacing ($\Delta_{H}=\Delta_{V}=\lambda/2$) placed $15$ m from the BS. User angles follow $\phi,\varphi\sim\mathcal{U}[-\pi/3,\pi/3]$. Near vs. far field is decided by the Fraunhofer distance $d_{\mathrm{f}}=2D^{2}/\lambda$, where $D$ is the largest aperture dimension. For the near field, $r\sim\mathcal{U}[d_{\mathrm{B}},\,d_{\mathrm{f}}/10]$ with $d_{\mathrm{B}}=2D$ the Björnson distance. For the far field $r\sim\mathcal{U}[d_{\mathrm{f}},\,5d_{\mathrm{f}}]$. Pilot SNR is taken as $10$ dB higher than the data SNR.
We benchmark against a conventional passive RIS under a fixed total power budget of $200$ mW. For the passive RIS, the entire budget is allocated for the BS. For the active RIS, $25\%$ of the budget powers the RIS internal amplifiers, and the BS radiates the remaining $75\%$, ensuring a fair comparison. Results are averaged over $1000$ independent channel & noise realizations.
With the adaptive pilot strategy and wide beams initialization technique for the Active RIS, under both near and far field regimes, the curve converges in around 7- 8 pilots.
Figs. 4 & 4 plot capacity bounds and pilot-aided achievable rates. For any pilot count, the active RIS yields higher rates, and its operational rate closely approaches its capacity bound, whereas the passive RIS remains well below. In the near field, the active RIS achieves $25.29$ bps/Hz vs. $14.31$ bps/Hz for the passive RIS. In the far field, the multiplicative fading effect can be seen much more clearly, where the active RIS attains nearly sixfold $21.01$ bps/Hz in comparison to $3.87$ bps/Hz for the passive RIS, under identical pilot configurations.
In the far field, 4 , the active RIS attains a final NMSE several dB below the minimum NMSE of the passive RIS, under both near and far-field approximation models. In the near field, 4 , the two architectures exhibit almost equivalent behavior under both models.
When the user is in the far field, both approximation models are nearly accurate. However, in the near field, the far-field approximation degrades. Due to not having any information regarding the distance, the far-field approximator can not reach its capacity.
For any fixed number of pilots, the active RIS consistently and markedly outperforms the passive RIS, requiring fewer training resources to achieve near-optimal performance, thereby reserving more channel uses for data and improving spectral efficiency.
Figure 5 shows the Capacity versus the power-splitting factor $\alpha=P_{\mathrm{RIS}}/P_{\mathrm{tot}}$ for $P_{\mathrm{tot}}\in\{200,500,1000\}$ mW. Increasing $\alpha$ initially improves the rate due to stronger RIS amplification, but the gain saturates, and drops as $\alpha\!\to\!1$ since the user transmit power $(1-\alpha)P_{\mathrm{tot}}$ vanishes, yielding an optimal split at intermediate-to-high $\alpha$.
Active RIS achieves better channel estimation and higher rates under the same total power budget, even with lower BS transmit power, since allocating power to RIS amplification mitigates double-path loss, making active RIS preferable to passive designs for reliable, high-efficiency links.
VI Conclusion
Equipping each RIS element with an amplifier fundamentally shifts the design trade-offs versus passive RIS: by turning multiplicative path loss into net link gain, active RIS can boost capacity and energy efficiency in 6G. Practical deployment, however, must confront hardware realities: amplifier nonlinearity, tight power budgets, and calibration drift. Semi-active designs that mix active and passive elements are also promising, e.g., sparse sensing with adaptive gain for scalable wideband estimation. Looking ahead, hybrid approaches that fuse model-based reasoning with data-driven learning can adapt to mobility and blockage while respecting thermal and power limits. Integrating codebooks, algorithms, and hardware co-design is likely to deliver meaningful spectral and energy efficiency gains as ARIS matures for next-generation networks.